Chapter 7, Section 2, Exercise 032 Find the expected count and the contribution to the chi-square statistic for the (B, E) cell in the two-way table below 36 70 21 34 93 39 37 69 29 G Total 35 58290 30 119 142 Total 127 166 135 123 551 Round your answer for the expected count to one decimal place, and your answer for the contribution to the chi-square statistic to three decimal places Expected count- contribution to the chi-square statistic-
Answer #1

Solution: Expected count for a cell is given as,

\text{Expected count}=\frac{(\text{sum of data in that row})\times(\text{sum of data in that column})}{\text{total data}}

S, for the (B,E) cell the expected count:

\text{Expected count}=\frac{(290)*(166)}{551} = 87.368

Contribution to the chi-square statistic for a cell is given by,

\text{Contribution to}\ \chi ^2=\frac{(\text{observed count - expected count})^2}{\text{expected count}}

\text{Contribution to}\ \chi ^2\text{ for (B,E)} = \frac{(93-87.368)^2}{87.368} = 0.363